An object projected upwards acquires a velocity of 9.8m/s when it reaches half of the maximum height. The maximum height reached is?

Grade:11

1 Answers

Arun
5 years ago
Assuming "projected upwards" means "projected VERTICALLY upwards". 
Then when falling from its max height (h) the object will achieve the same downward velocity = 9.8 m/s as it had when vertically rising at h/2. 
V = √2gH = √2g(h/2) = 9.8 m/s 
(9.8)² = 2g(h/2) 
h/2 = (9.8)²/2(9.8) = 4.9
h = 2(4.9) = 9.8 m